2. Incremental costs - Initial and terminal cash flow

Consider the case of Marston Manufacturing Company:

Marston Manufacturing Company is considering a project that requires an investment in new equipment of $3,780,000. Under the new tax law, the equipment is eligible for 100% bonus depreciation at t = 0 so the equipment will be fully depreciated at the time of purchase. Marston estimates that its accounts receivable and inventories need to increase by $720,000 to support the new project, some of which is financed by a $288,000 increase in spontaneous liabilities (accounts payable and accruals). The company's tax rate is 25%.

The after-tax cost of Marston’s new equipment is

A. $2,835,000

B. $288,000

C. 3,600,000

Marston’s initial net investment outlay is

A. $2,979,000

B. $3,267,000

C. $2,835,000

Suppose Marston’s new equipment is expected to sell for $600,000 at the end of its four-year useful life, and at the same time, the firm expects to recover all of its net operating working capital (NOWC) investment. Remember, that under the new tax law, this equipment was fully depreciated at t = 0. If the firm’s tax rate is 25%, what is the project’s total termination cash flow?

A. $582,000

B. $450,000

C. $882,000

D. $600,000
Answer #1

A). After-tax cost of new equipment = cost*(1-Tax rate) = 3,780,000*(1-25%) = 2,835,000 (Option A)

B). Initial investment = equipment cost + increase in NOWC - (depreciation*Tax rate)

= 3,780,000 + (720,000 - 288,000) - (3,780,000*25%) = 3,267,000 (Option B)

C). Termination cash flow = after-tax salvage value of equipment + recovery of NOWC

= 600,000*(1-25%) + 432,000 = 882,000 (Option C)